The perimeter of a rectangular garden is 40yd, and its area is 36yd^2. How many yards long is the longest side of the garden? (yd^2 = Square Feet)

Answer by checkley79(3341)

P=2L+2W
A=LW
36=LW
L=36/W
40=2(36/W)+2W
40=72/W+2W
40=(72+2W*W)/W
40=(72+2W^2)/W
40W=2W^2+72
2W^2-40W+72=0
2(W^2-20W+36)=0
2(W-18)(W-2)=0
W-18=0
W=18 ANS.
L=36/18=2 ANS.
W-2=0
W=2 ANS.
################
L=36/2
L=18 ANS.